Need help with the following:
Solve the inequality |3x - 1| < 7 and write the answer in the set notation.

Answer by venugopalramana(3286)

Need help with the following:
Solve the inequality |3x - 1| < 7 and write the answer in the set notation.
CASE 1..
3X-1 >0
|3X-1|=3X-1<7
3X<7+1=8
X<8/3
CASE 2.
3X-1<0
|3X-1|=-(3X-1)=-3X+1<7
-3X<7-1=6
-X<6/3=2
X>2
SO SOLUTION SET IS {X:X>2 AND X<8/3}